\int \:\frac{x^2+1}{x^2}dx
1
Expert's answer
2020-09-14T18:10:25-0400
"\int\dfrac{x^2+1}{x^2}dx=\int dx+\int\dfrac{1}{x^2}dx=x-\dfrac{1}{x}+C"